Search
Text Size:
Smaller Text Normal Text Larger Text

Facebook page YouTube channel

BlueSky profile YouTube channel
P68-030

P68-030: Oscillating wildly - Characterising the nature of unusually pulsating stars II

P.I.: V. Antoci

This is a Monitoring program using FIES.

Instrument setup

  • FIES

  • High-res fiber F4

Afternoon instructions

  • Run afternoon calibrations: fies-calibs -fib 4

Observing instructions

  • On the Queue page, make sure the OB-generated observing script is written to disk, or else 'compile'

  • Go to directory: cd ~/obs/scripts/68-030/ and do ls -lrt

  • Start the requested OB generated script


Program Tracking
Proposal 68-030
Semester P68: 20231001 - 20240401
Principal Investigator Victoria Antoci
Programme Oscillating wildly-Characterising the nature of unusually pulsating stars II
Instrument FIES
Time Allocation 2 separate observations of 10 targets
Total Nights 10.6h
Type Mon
Notes Fixed dates to be scheduled
Total Time Spent 08:14 (hh:mm)
Proposal Date Time Type Targets Files Comments Status
68-030 2023-10-08 00:00 Mon thick clouds all night Failed
68-030 2023-10-14 00:22 Mon HD187258 FIGj140094 - FIGj140096 Observed
68-030 2023-10-19 00:20 Mon HD187258 FIGj190099 - FIGj190101 Target HD187258 observed with clear skies, high wind, seeing of around 1.5 arcsec. Target HD201032 could not be observed because of humidity out of limits. Observed
68-030 2023-11-04 00:35 Mon HD27230 FIGk040080 - FIGk040082 Clear skies. Seeing over 1 arcsec. Observed
68-030 2023-11-07 00:35 Mon HD27230 FIGk070116 - FIGk070118 Observed during clear skies. Dimm seeing around 1.2 arcsec, but following the fussiness of the guide star my guess is, that the seeing is worse than this. Observed
68-030 2023-11-08 00:35 Mon HD2523 FIGk080066 - FIGk080068 Target observed with clear skies, seeing of 1.0". Strong winds were present, but did not disturb the observation. Observed
68-030 2023-11-30 00:00 Mon No observations due to weather. Failed
68-030 2024-02-13 00:34 Mon HD112515 FIHb130001 - FIHb130003 Observed with thin clouds and seeing around 1 arcsec. S/N estimates from exposure meter were 155 (O9), 165 (F3), 212 (G7) and 295 (M2). Observed
68-030 2024-02-15 00:41 Mon HD86167 FIHb150074 - FIHb150076 Observed with clear skies and seeing around 0.8 arcseconds. Maybe some errors during calibration, please check quality of calibration frames. Observed
68-030 2024-02-19 00:40 Mon HD112515 FIHb190105 - FIHb190107 Target observed through thin clouds and seeing of around 1.5 arcsec. Added 400s to the exposure because of the clouds, in order to have a higher SNR. Observed
68-030 2024-02-20 00:33 Mon HD86167 FIHb200101 - FIHb200103 Observed in clear skies with seeing around 1.0" Observed
68-030 2024-03-09 01:14 Mon HD104957, HD113697 FIHc090099 - FIHc090101, FIHc090102 - FIHc090104 All targets observed with clear skies and seeing between 0.8" and 0.9" Observed
68-030 2024-03-14 00:32 Mon HD113697 FIHc140004 - FIHc140006 Clear skies, and a seeing of around 0.5 arcsec. We had some problems with the fiber guider, so we had to do a couple CTRL-C, but it worked out in the end! FIHc140004 - FIHc140006 were successful observations Observed
68-030 2024-03-17 01:13 Mon HD114839, HD104957 FIHc170100 - FIHc170105 All targets observed with clear skies within seeing constraints. Observed
68-030 2024-03-21 00:00 Mon Observation was missed due to the telescope being closed because of harsh weather conditions. Failed
68-030 2024-03-25 00:00 Mon - - The target (HD158251) was missed due to bad weather. Failed
68-030 2024-03-30 00:20 Mon HD158251 FIHc300120 - FIHc300122 Observed in thin clouds with seeing around 1.5" Observed

 

The above information is collected from the ToO/Mon reports that have been created by the observer.

Should any of the observations have been carried out using the OB system, you will find additional information HERE

Back to top Last modified: September 27 2023